OpenStudy (anonymous):

double integral of (x cos xy dydx) from pi/2 to pi and from 1 to 2

sam (.sam.):

\[\int\limits \int\limits xcos(xy) ~dydx\] Do the one in bracket first \[\int\limits\limits[ \int\limits\limits xcos(xy) ~dy]dx\]

OpenStudy (agent0smith):

\[\Large \int\limits_{1}^{2} \left[ \int\limits_{\pi/2}^{\pi} x \cos (xy) ~dy \right] dx \]as sam said, do the part in square brackets first. Integrate that w.r.t. y, while treating x as if it's a constant.